# Effective Hamiltonian approach to the Quantum Phase transitions in the extended Jaynes-Cummings model

### 11 Dec 2023 v2

@@ II.2 Complex discretization approximation
as an evaluation. This is because $H^{\dagger}_{\text{dis}}$ would creat a gain effect since its complex eigenvalues have a positive imaginary part, which is expected to reduce the errors. To justify this proposition, we reexamine the spin dynamics of $s=1$ by $\widetilde{H}_{\text{dis}}$. As shown in Fig. 3, $\widetilde{H}_{\text{dis}}$ provides a reliable assessment of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ for $\eta=0.05$ and $0.1$, with a significant reduction of error by at least a factor of ten. The explicit illustration for the evolution can be found in Fig. A2 in Appendix I
@@
Figure 5: (Color online) The evolution of survival probability is plotted for selected $s$ and $\eta$. For panels (a) and (b), the evolution is obtained by $H_{\text{dis}}$. While for panel (c), $\widetilde{H}_{\text{dis}}$ is used to obtain convergent results for $\eta=0.05,0.1$. As for $\eta=0.5$, the evolution is determined by the combination of $H_{\text{dis}}$ and $\widetilde{H}_{\text{dis}}$, as stated at the end of II.2. The Roman numbers denote the different regions in Fig. 4, to which the chosen values of $s$ and $\eta$ belong. For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
@@ III Quantum Phase Transition in the single-excitation subspace
Actually, the CDA method transforms the open quantum system into a closed many-body system, allowing for the conventional approach to quantum phase transition sachdev to be applied. An overview to the main result of this section is provided first. After then, the spectrum of $H_{\text{dis}}$ is studied, with a focus on the energy gap of ground state, which can be used to identify the quantum phase transition. Finally, the ground state function is analyzed in details, which shows a significant change in the population of excitation with an increase of the coupling strength $\eta$ and thus indicates an intrinsic transition of spin dynamics.
@@
The main result is summarized as the phase diagram shown in Fig. 4. Evidently, phase diagram can ge divided into three regions, labelled as I, II and III. The region I denotes the dissipative phase, where the spin dynamics decay rapidly and excitation is absorbed into environment. In contrast, the region II denotes the localized phase, where excitation may be localized in spin system by a finite probability. Between this two regions, a third region, named as Phase III, can be found for small $s$. In this region, the spin dynamics shows an intermediate behavior between the fast decaying and localization. Our calculation shows that the dynamics of system in this region displays an stretched-like feature. This feature is a result of the competition of the dissipation induced by coupling to environment and the localization from the opening of energy gap of ground state. The boundary between region I and II or III is decided by critical value $\eta_{I}$ or $\eta_{II}$. At $\eta_{I}$, the energy gap of ground state opens, while at $\eta_{II}$ the population of excitation of the ground state becomes maximal in the spin degree of freedom. For large $s$, $\eta_{I}$ and $\eta_{II}$ merge into a single one, at which the spin dynamics shows an intermediate feature. The relevance of $\eta_{I}$ and $\eta_{II}$ to $N$ and $R$ is studied in Appendix II
The survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ with $|\psi(0)\rangle=|1\rangle|0\rangle^{\otimes k}$ is evaluated to illustrate the distinct dynamics in three regions. As shown in Fig. 5, dependent on the values of $s$ and $\eta$ belonging to different regions in Fig. 4, $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ can display fast decaying and localization respectively. In order to display the intermediate dynamics in region III, a numerical fitting for the long-term evolution of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ is provided in Fig. A4 in Appendix II, based on the fitting function $B\exp\left(-At^{\beta}\right)$. It is found that the exponential index $\beta$ has the magnitude of order $10^{-1}$, which clearly demonstrate the stretched-like feature of the evolution.
@@ III.2 The spectrum of $H_{\text{dis}}$
Figure 6: (Color online)Plots of the spectrum of $H_{\text{dis}}$ for different $s$ and $\eta$. For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
To understand the phase diagram, the spectrum of $H_{\text{dis}}$ is studied in this subsection, obtained by exactly diagonalizing $H_{\text{dis}}$. As shown in Fig. 6, the eigenvalues exhibit a non-positive imaginary part, characterizing the dissipative spin dynamics. Furthermore, their real parts are confined at interval $\left(0,120\right)$. This is a consequence of CDA approach, that the continuum of environment may be discretized as a finite spectrum bound at interval $\left(0,2R\omega_{c}\right)$.
As for the non-Hermitianity of $H_{\text{dis}}$, the ground state can be defined as the right eigenfunction with the minimal real part of its eigenvalue. For small $\eta$, the ground state would correspond to the energy level with the real part close to zero, as shown by left column in Fig.6. However when $\eta$ increases beyond a critical value $\eta_{I}$ related to $s$, a single level emerges in the region $\Re[E]<0$, as shown by the middle and right columns in Fig.6. In this case, the ground state becomes separated from the level band by a energy gap, which is equal to the real part of its eigenvalue. In addition, the gap increases with increment of $\eta$.
Physically, the single ground state corresponds to the photon-bound state, which has been known to be associated with the localized dynamics of system kofman1994 ; bs . Nevertheless, as shown in Fig. 5, the opening of energy gap does not guarantee the emergence of localization in the dynamics, a stretched-like decaying observed when $\eta>\eta_{I}$. Additionally, localized dynamics can only be observed when $\eta$ surpasses a second critical value. Consequently, to understand the intricate behaviors, it is necessary to inspect the ground state function.
@@ III.3 The ground state function
The ground state function can be obtained directly by solving $H_{\text{dis}}$, which can be written formally as $|g\rangle=\alpha|e\rangle|0\rangle+|g\rangle\left(\sum_{k}\beta_{k}(t)|1\rangle_{k}\right)$. Here, $|e\rangle|0\rangle\equiv|1\rangle$ denotes excitation in the spin, while $|g\rangle|1\rangle_{k}\equiv|k\rangle(k=2,3,\cdots,N+1)$ denotes the excitation in the $k$-th mode of environment. The population of excitation on $|1\rangle$ has a significant influence on the localized evolution of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$. In Fig.7, the position of maximal population (PMP) on the basis state $|k\rangle$ is explored for the ground state when $s=0,0.2$ and $1$. For the three situations, the energy gap of ground state opens at critical point $\eta_{I}=0.0076,0.02335$ and $0.1$ respectively, as illustrated in Fig.4. However, as shown in Fig. 7 (a) and (b), PMP at $|1\rangle$ can happen only when $\eta$ is beyond a critical value, labeled as $\eta_{II}$. This can explain the evolution illustrated in Fig. 5(a) and (b), in which the localized behavior of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ can occur only for $\eta$ considerably greater than $\eta_{I}$. Thus, the stretched-like dynamics is a result of the competition between the energy gap, which protects excitation against the spontaneous emission into environment, and the intrinsic dissipation in the ground state function, which does not support PMP at $|1\rangle$.
@@
For $s=1$, $\eta_{II}$ is found to be almost equal to $\eta_{I}$. Thus, the stretched-like evolution can be found at critical point $\eta_{I}=0.1$. Additionally, Fig. A5 (c) in Appendix II reveals that $\eta_{II}$ is smaller than $\eta_{I}$. This difference, however, tends to diminish when $N\rightarrow\infty$. Similar picture can also be found for other values of $s$.
@@
In addition, the average value $\langle\sigma_{z}\rangle$ for the ground state is evaluated in order to detect phase transition. As shown by Fig. 5, $\langle\sigma_{z}\rangle$ remains continuous for $s=0,0.2$, yet a jump is observed at $\eta=0.1$ for $s=1$, implying that the transition for $s=1$ may be distinct from $s=0,0.1$.
@@
Conclusively, we have demonstrated the presence of a second critical point $\eta_{II}$ by analyzing the ground state function. For small $s$, $\eta_{II}$ is distinct from $\eta_{I}$, and an intermediate region can be identified where the spin dynamics shows stretched-like behavior. As $s$ increases, $\eta_{II}$ tends to be equal to $\eta_{I}$, and the stretched-like spin dynamics can happen at critical point. Furthermore, the two scenarios can be distinguished by evaluating $\langle\sigma_{z}\rangle$ for the ground state. If $\eta_{II}$ not equal to $\eta_{I}$, $\langle\sigma_{z}\rangle$ is continuous at $\eta_{II}$. Conversely,If $\eta_{II}$ tends to be equal to $\eta_{I}$, $\langle\sigma_{z}\rangle$ exhibits a jump c at $\eta_{II}$.
@@
Figure 8: (Color online) The comparative plots of spin dynamics in the double-excitation subspace (empty circle) and single-excitation subspace (solid line). The spin dynamics in single-excitation subspace is evaluate by $H_{\text{dis}}$ for $s=0,0.2$ and $\widetilde{H}_{\text{dis}}$ for $s=1$. For the calculation in double-excitation subspace, $\widetilde{H}_{\text{dis}}$ is solved by choosing $N=200$ in order to obtain the convergent result. For all plots, $R=6,\Delta=1,\omega_{c}=10$ are chosen.
@@
Tthe wave function in the double-excitation subspace can be expressed as
@@
Evidently, $|e\rangle$ is correlated to the status of environment. To characterize the spin dynamics,
@@
is evaluated for arbitrary time $t$, which denots the probability of the single excitation saved by the spin. The initial state is chosen as $|\phi(0)\rangle=1/\sqrt{N}|e\rangle\sum_{k=1}^{N}a^{\dagger}_{k}|0\rangle^{\otimes N}$ to avoid the dependence on special mode in environment. $N=200$ is chosen for the simulation, which corresponds to the Hilbert space of dimension $\sim 2.0\times 10^{5}$. For larger $N$, the calculation becomes too consumptive to implement. To offset the recurrence induced by finite $N$, $\widetilde{H}_{\text{dis}}$ is used to find convergent $P_{e}(t)$ in this evaluation, which is illustrated by empty circles in Fig. 8. Because of finite $N$, the calculation becomes fluctuated when $t>\sim 10$. For $s=1$, the calculation shows a slow convergency, which induces strong fluctuation and large computational error.
It is observed that $P_{e}(t)$ displays similar evolution as $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ in single excitation subspace. This has been verified further by Gauss quadratures method in real space discretization , as illustrated by Fig. A6 in Appendix III. This observation would be attributed to the fact that the discretized environment corresponds to an interaction-free bosonic system, in which the excitations can be populated freely at the mode of environment without any interference. Thus, the decay of excitation in spin would be independent of the status of environment. This implies that the phase diagram in single-excitation subspace would be robust against the increase of energy in the total system.
We try to find explanation of spin dynamics by analyzing spectrum of $H_{\text{dis}}$. However, the spectrum of $H_{\text{dis}}$ does not provide sufficient evidence to explain the spin dynamics in terms of the ground state, as is possible in the single-excitation subspace. Calculations show that the energy gap can remain finite even when $\eta<\eta_{I}$, and the ground state can have positive energy in such cases. Furthermore, the number of eigenvalues with negative real parts depends on the value of $\eta$. Consequently, the bound state is absent in the multiple-excitation subspace, and the dynamics of spin cannot be explained from the single-excitation perspective.

### 06 Apr 2024 v3

@@ II.3 Introduction of $\tilde{H}_{\text{dis}}$
as an evaluation. In practice, $H^{\dagger}_{\text{dis}}$ can create a gain effect since its complex eigenvalues have a positive imaginary part. Thus, it is expected to reduce the unwanted decreasing induced by finite $N$. To justify this proposition, we reexamine the spin dynamics of $s=1$ by $\widetilde{H}_{\text{dis}}$. As shown in Fig. 3(a), $\widetilde{H}_{\text{dis}}$ provides a reliable assessment of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ for $\eta=0.05$ and $0.1$, with a significant reduction of error by at least a factor of 10. The explicit illustration for the evolution can be found in Fig. A2 in Appendix I
@@
To justify the introduction of $\tilde{H}_{\text{dis}}$, we offer the analytical evidence in this section as an addition to the numerical evidence presented in Fig. A3 in Appendix I. The objective is to ascertain the long-term behavior of $\alpha(t)$, which can be determined analytically by solving Eq. (5). It is worth noting that the enduring features of $\alpha(t)$ are determined by the bound state kofman1994 , which occurs when
$\Delta-\int_{0}^{\infty}\text{d}\omega\frac{J(\omega)}{\omega}<0.$ (23)
For the Ohmic-type environment defined by Eq. (2), one gets
$\eta>\frac{\Delta}{\omega_{c}\Gamma(s)}:=\eta_{c}$ (24)
Evidently, $\eta_{c}=0.1$ when $s=1$ and $\Delta=1,\omega_{c}=10$. By solving Eq.(6), a single bound state can be found at $E_{b}=-2.369$ when $\eta=0.5$. At $E_{b}$, the Laplace transformation $\alpha(p)$ defined in Eq. (5), can be obtained readily by finding $\lim_{E\rightarrow E_{b}}\left(E-E_{b}\right)\alpha(p=\mathbbm{i}E)$. This leads to the approximation of $\alpha(t)\approx-0.739894e^{-\mathbbm{i}2.369t}$ and a resulting survival probability of $\sim 0.547439$. In contrast, the calculation by $\tilde{H}_{\text{dis}}$ shows a slight oscillation around $\sim 0.54744$ as depicted in Fig. 3(b). This perfect agreement between the analytical and numerical approaches shows unambiguously the ability of $\tilde{H}_{\text{dis}}$ to predict the steady feature of $\alpha(t)$.
In the case of no bound state, the integral in Eq. (6) becomes divergent, leading to $\alpha(t)$ approaching zero and the survival probability decreasing as $t$ increases. Due to the lack of an analytical expression for $\alpha(t)$, we have to rely on numerical methods. The calculations using $\tilde{H}_{\text{dis}}$ in Fig. A2 (a) and A2(b) show a reasonably prediction for the steady behavior of survival probability, based on the trend with increasing $N$.
@@
Figure 5: (Color online) The evolution of survival probability with time $t$ (in units of $1/\Delta$) is plotted for selected $s$ and $\eta$. For panels (a) and (b), the evolution is obtained by $H_{\text{dis}}$. While for panel (c), $\widetilde{H}_{\text{dis}}$ is used to obtain convergent results for $\eta=0.05,0.1$. As for $\eta=0.5$, the evolution is determined by the combination of $H_{\text{dis}}$ and $\widetilde{H}_{\text{dis}}$, as stated at the end of Sec. II.2. The Roman numbers denote the different regions in Fig. 4, to which the chosen values of $s$ and $\eta$ belong. For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
@@ III Quantum Phase Transition in the single-excitation subspace
Actually, the CDA method transforms the open quantum system into a closed many-body system, allowing for the conventional approach to quantum phase transition sachdev to be applied. An overview to the main result of this section is provided first. After that, the spectrum of $H_{\text{dis}}$ is studied, with a focus on the energy gap of A2 ground state, which can be used to identify the quantum phase transition. Finally, the ground state function is analyzed in details, which shows a significant change in the population of excitation with an increase of the coupling strength $\eta$ and thus indicates an intrinsic transition of spin dynamics.
@@
The main result is summarized as the phase diagram shown in Fig. 4. Evidently, the phase diagram can ge divided into three regions, labeled as I, II and III. Region I denotes the dissipative phase, where the spin dynamics decay rapidly and excitation is absorbed into the environment. In contrast, region II denotes the localized phase, where excitation may be localized in the spin system by a finite probability. Between these two regions, a third region, named as phase III, can be found for small $s$. In this region, the spin dynamics shows an intermediate behavior between the fast decaying and localization. Our calculation shows that the dynamics of the system in this region displays an stretched-like feature. This feature is a result of the competition of the dissipation induced by coupling to the environment and the localization from the opening of energy gap of the ground state. The boundary between region I and II or III is decided by the critical value $\eta_{I}$ or $\eta_{II}$. At $\eta_{I}$, the energy gap of ground state opens, while at $\eta_{II}$ the population of excitation of the ground state becomes maximal in the spin degree of freedom. For large $s$, $\eta_{I}$ and $\eta_{II}$ merge into a single one, at which the spin dynamics shows an intermediate feature. The relevance of $\eta_{I}$ and $\eta_{II}$ to $N$ and $R$ is studied in Appendix II
The survival probability $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ with $|\psi(0)\rangle=|1\rangle|0\rangle^{\otimes k}$ is evaluated to illustrate the distinct dynamics in three regions. As shown in Fig. 5, dependent on the values of $s$ and $\eta$ belonging to different regions in Fig. 4, $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ can display fast decaying and localization respectively. To display the intermediate dynamics in region III, a numerical fitting for the long-term evolution of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ is provided in Fig. A4 in Appendix II, based on the fitting function $B\exp\left(-At^{\beta}\right)$. It is found that the exponential index $\beta$ has the magnitude of order $10^{-1}$, which clearly demonstrates the stretched-like feature of the evolution.
@@ III.2 The spectrum of $H_{\text{dis}}$
Figure 6: (Color online)Plots of the spectrum (in units of $\Delta$) of $H_{\text{dis}}$ for different $s$ and $\eta$. For all plots, $N=2000,R=6,\Delta=1$, and $\omega_{c}=10$ are chosen.
To understand the phase diagram, the spectrum of $H_{\text{dis}}$ is studied in this subsection, obtained by exactly diagonalizing $H_{\text{dis}}$. As shown in Fig. 6, the eigenvalues exhibit a non-positive imaginary part, characterizing the dissipative spin dynamics. Furthermore, their real parts are confined at interval $\left(0,120\right)$. This is a consequence of the CDA approach, where the continuum of the environment may be discretized as a finite spectrum bound at interval $\left(0,2R\omega_{c}\right)$.
As for the non-Hermitianity of $H_{\text{dis}}$, the ground state can be defined as the right eigenfunction with the minimal real part of its eigenvalue. For small $\eta$, the ground state would correspond to the energy level with the real part close to zero, as shown by the left column in Fig.6. However when $\eta$ increases beyond a critical value $\eta_{I}$ related to $s$, a single level emerges in the region $\text{Re}[E]<0$, as shown by the middle and right columns in Fig.6. In this case, the ground state becomes separated from the level band by a energy gap, which is equal to the real part of its eigenvalue. In addition, the gap increases with increment of $\eta$.
Physically, the single ground state corresponds to the photon-bound state, which has been known to be associated with the localized dynamics of a system kofman1994 ; bs . Nevertheless, as shown in Fig. 5, the opening of the energy gap does not guarantee the emergence of localization in the dynamics, a stretched-like decaying observed when $\eta>\eta_{I}$. Additionally, localized dynamics can only be observed when $\eta$ surpasses a second critical value. Consequently, to understand the intricate behaviors, it is necessary to inspect the ground state function.
@@ III.3 The ground state function
The ground state function can be obtained directly by solving $H_{\text{dis}}$, which can be written formally as $|g\rangle=\alpha|e\rangle|0\rangle+|g\rangle\left(\sum_{k}\beta_{k}(t)|1\rangle_{k}\right)$. Here, $|e\rangle|0\rangle\equiv|1\rangle$ denotes excitation in the spin, while $|g\rangle|1\rangle_{k}\equiv|k\rangle(k=2,3,\cdots,N+1)$ denotes the excitation in the $k$-th mode of the environment. The population of excitation on $|1\rangle$ has a significant influence on the localized evolution of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$. In Fig. 7, the position of maximal population (PMP) on the basis state $|k\rangle$ is explored for the ground state when $s=0,0.2$ and $1$. For the three situations, the energy gap of ground state opens at critical point $\eta_{I}=0.0076,0.02335$ and $0.1$ respectively, as illustrated in Fig.4. However, as shown in Fig. 7 (a) and (b), PMP at $|1\rangle$ can happen only when $\eta$ is beyond a critical value, labeled as $\eta_{II}$. This can explain the evolution illustrated in Fig. 5(a) and 5(b), in which the localized behavior of $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ can occur only for $\eta$ considerably greater than $\eta_{I}$. Thus, the stretched-like dynamics is a result of the competition between the energy gap, which protects excitation against the spontaneous emission into the environment, and the intrinsic dissipation in the ground state function, which does not support PMP at $|1\rangle$.
@@
For $s=1$, $\eta_{II}$ is found to be almost equal to $\eta_{I}$. Thus, the stretched-like evolution can be found at critical point $\eta_{I}=0.1$. Additionally, Fig. A5 (c) in Appendix II reveals that $\eta_{II}$ is smaller than $\eta_{I}$. This difference, however, tends to diminish when $N\rightarrow\infty$. A similar picture can also be found for other values of $s$.
@@
In addition, the average value $\langle\sigma_{z}\rangle$ for the ground state is evaluated to detect phase transition. As shown by Fig. 5, $\langle\sigma_{z}\rangle$ remains continuous for $s=0,0.2$, yet a jump is observed at $\eta=0.1$ for $s=1$, implying that the transition for $s=1$ may be distinct from $s=0,0.1$.
@@
Conclusively, we demonstrated the presence of a second critical point $\eta_{II}$ by analyzing the ground state function. For small $s$, $\eta_{II}$ is distinct from $\eta_{I}$, and an intermediate region can be identified where the spin dynamics shows stretched-like behavior. As $s$ increases, $\eta_{II}$ tends to be equal to $\eta_{I}$, and the stretched-like spin dynamics can happen at a critical point. Furthermore, the two scenarios can be distinguished by evaluating $\langle\sigma_{z}\rangle$ for the ground state. If $\eta_{II}$ does not equal to $\eta_{I}$, $\langle\sigma_{z}\rangle$ is continuous at $\eta_{II}$. Conversely, if $\eta_{II}$ tends to be equal to $\eta_{I}$, $\langle\sigma_{z}\rangle$ exhibits a jump c at $\eta_{II}$.
@@
Figure 8: (Color online) The comparative plots of spin dynamics in the double-excitation subspace (empty circle) and single-excitation subspace (solid line). The spin dynamics in single-excitation subspace is evaluate by $H_{\text{dis}}$ for $s=0,0.2$ and $\widetilde{H}_{\text{dis}}$ for $s=1$. For the calculation in double-excitation subspace, $\widetilde{H}_{\text{dis}}$ is solved by choosing $N=200$ to obtain the convergent result. For all plots, $R=6,\Delta=1,\omega_{c}=10$ are chosen. The evolution time $t$ is in units of $1/\Delta$.
@@
The wave function in the double-excitation subspace can be expressed as
@@
Evidently, $|e\rangle$ is correlated to the status of the environment. To characterize the spin dynamics,
@@
is evaluated for arbitrary time $t$, which denotes the probability of the single excitation saved by the spin. The initial state is chosen as $|\phi(0)\rangle=1/\sqrt{N}|e\rangle\sum_{k=1}^{N}a^{\dagger}_{k}|0\rangle^{\otimes N}$ to avoid the dependence on the special mode in the environment. $N=200$ is chosen for the simulation, which corresponds to the Hilbert space of dimension $\sim 2.0\times 10^{5}$. For larger $N$, the calculation becomes too consumptive to implement. To offset the recurrence induced by finite $N$, $\widetilde{H}_{\text{dis}}$ is used to find convergent $P_{e}(t)$ in this evaluation, which is illustrated by the empty circles in Fig. 8. Because of finite $N$, the calculation becomes fluctuated when $t>\sim 10$. For $s=1$, the calculation shows a slow convergency, which induces strong fluctuation and large computational error.
It is observed that $P_{e}(t)$ displays a similar evolution as $\left|\langle\psi(t)|\psi(0)\rangle\right|^{2}$ in the single excitation subspace. This has been verified further by the Gauss quadratures method in real space discretization , as illustrated by Fig. A6 in Appendix III. This observation would be attributed to the fact that the discretized environment corresponds to an interaction-free bosonic system, in which the excitations can be populated freely at the mode of the environment without any interference. Thus, the decay of excitation in spin would be independent of the status of the environment. This implies that the phase diagram in single-excitation subspace would be robust against the increase of energy in the total system.
We try to find an explanation of spin dynamics by analyzing the spectrum of $H_{\text{dis}}$. However, the spectrum of $H_{\text{dis}}$ does not provide sufficient evidence to explain the spin dynamics in terms of the ground state, as is possible in the single-excitation subspace. Calculations show that the energy gap can remain finite even when $\eta<\eta_{I}$, and the ground state can have positive energy in such cases. Furthermore, the number of eigenvalues with negative real parts depends on the value of $\eta$. Consequently, the bound state is absent in the multiple-excitation subspace, and the dynamics of spin cannot be explained from the single-excitation perspective.